Circular No. 6362

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 B2 (HYAKUTAKE) D. Lis, J. Keene, K. Young, and T. Phillips, California Institute of

Technology Submillimeter Observatory; E. Bergin, Smithsonian Astrophysical Observatory; P. Goldsmith, Cornell University; D. Bockelée-Morvan, J. Crovisier, and D. Gautier, Observatoire de Paris-Meudon; A. Wootten, National Radio Astronomy Observatory; D. Despois, Observatoire de Bordeaux; and T. Owen, University of Hawaii, report the detection of the 110-000 rotational line of HDO at 464.9 GHz in comet C/1996 B2, using the Caltech Submillimeter Observatory on Mar. 23.5 and 24.5 UT. The integrated line intensity on the antenna temperature scale is 0.34 ± 0.05 K km/s. Assuming local thermal equilibrium at 80 K, we derive Q(HDO)= 1.0 × 10^{26} mol/s. This corresponds to a [HDO]/[H_{2}O] ratio of 2.8 × 10^{−4}, using the water-production rate measured from OH observations with IUE (IAUC 6355). These preliminary results suggest a [D]/[H] ratio of 1.4 × 10^{−4}, which is by a factor of two lower than that measured by mass spectroscopy in water in comet 1P/Halley (Balsiger et al. 1995, J. Geophys. Res. 100, 5827; Eberhardt et al. 1995, A.Ap. 302, 301) and very close to the standard mean ocean water value of 1.6 × 10^{−4}.”

A. Wootten and B. Butler, National Radio Astronomy Observatory (NRAO); D. Bockelée-Morvan and J. Crovisier, Observatoire de Paris-Meudon; P. Palmer, University of Chicago; D. Despois, Observatoire de Bordeaux; and D. Yeomans, Jet Propulsion Laboratory, report the detection of the (J,K) = (1,1) and (3,3) inversion-rotation lines of ammonia at 1.3 cm on Mar. 24.5 UT, using the NRAO 43-m telescope at Green Bank: “We measure an integrated intensity in the (3,3) line of 0.22 ± 0.013 and 0.13 ± 0.015 K km s^{−1} in the main component and inner hyperfine components of the (1,1) line. From these four lines, we derive the rotational temperature of the ammonia molecules to be 60 K, assuming an abundance ratio of ortho- to para-ammonia of 3. Ammonia appears to be produced at a rate Q(NH_{3}) = 1 × 10^{27} mol/s, or only 0.25 percent that of water molecules, as reported on IAUC 6355 and derived from IUE observations of OH. Such an abundance lies at the extreme lower range of values deduced from 1P/Halley observations of NH and NH_{2} radicals, and much below those measured for 1P/Halley from Giotto mass spectroscopy.”

1996 March 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green